Express sin 5θ and cos 5θ/ cos θ in terms of sin θ
1
Expert's answer
2020-05-25T20:48:16-0400

"\sin(5\theta) = \sin(2\theta+3\theta)=\sin(2\theta)\cos(3\theta)+\sin(3\theta)cos(2\theta)\newline
\sin(3\theta) = 3\sin(\theta)\cos^2(\theta) – \sin^3(\theta)\newline
\cos(3\theta) = \cos^3(\theta) – 3\cos(\theta)\sin^2(\theta)\newline
\sin(2\theta) = 2\sin(\theta)\cos(\theta)\newline
\cos(2\theta)=1-\sin^2(\theta)\newline
\sin(5\theta)= 2\sin(\theta)\cos(\theta) \times (\cos^3(\theta) – 3\cos(\theta)\sin^2(\theta)) +\newline
+(3\sin(\theta)\cos^2(\theta) – \sin^3(\theta))\times(1-\sin^2(\theta))=\newline
=2\sin(\theta)\cos^4(\theta)-6\sin^3(\theta)\cos^2(\theta)+3\sin(\theta)\cos^2(\theta)-\newline
-3\sin^3(\theta)cos^2(\theta)-\sin^3(\theta)+\sin^5(\theta)=\newline
=2\sin(\theta)(1-\sin^2(\theta))^2-6\sin^3(\theta)(1-\sin^2(\theta))+3\sin(\theta)(1-\sin^2(\theta))-\newline
-3\sin^3(\theta)(1-\sin^2(\theta))-\sin^3(\theta)+\sin^5(\theta)=\newline
=2\sin(\theta)-4\sin^3(\theta)+2\sin^5(\theta)-6\sin^3(\theta)+6\sin^5(\theta)+3\sin(\theta)-3\sin^3(\theta)-\newline
-3\sin^3(\theta)+3\sin^5(\theta)-\sin^3(\theta)+\sin^5(\theta)=\newline
=13\sin^5(\theta)-17\sin^3(\theta)+5\sin(\theta)\newline"


"\cos(5\theta)=\cos(3\theta+2\theta)=\cos(3\theta)\cos(2\theta)-\sin(3\theta)\sin(2\theta)"

From the previous paragraph:

"\sin(3\theta) = 3\sin(\theta)\cos^2(\theta) – \sin^3(\theta)\newline
\cos(3\theta) = \cos^3(\theta) – 3\cos(\theta)\sin^2(\theta)\newline
\sin(2\theta) = 2\sin(\theta)\cos(\theta)\newline
\cos(2\theta)=1-\sin^2(\theta)\newline"

So:

"\cos(5\theta)= (\cos^3(\theta) – 3\cos(\theta)\sin^2(\theta))(1-\sin^2(\theta))-\newline
- (3\sin(\theta)\cos^2(\theta) – \sin^3(\theta))\times2\sin(\theta)\cos(\theta))=\newline
=\cos(\theta)(\cos^2(\theta) – 3\sin^2(\theta))(1-\sin^2(\theta))-\newline
-2\sin(\theta)\cos(\theta)(3\sin(\theta)\cos^2(\theta) – \sin^3(\theta))=\newline
=\cos(\theta)((1-4\sin^2(\theta))(1-\sin^2(\theta))-2\sin(\theta)(3\sin(\theta)-4\sin^3(\theta))=\newline
=\cos(\theta)(1-5\sin^2(\theta)+4\sin^4(\theta)-6\sin^2(\theta)+8\sin^4(\theta))=\newline
=\cos(\theta)(1-11\sin^2(\theta)+12\sin^4(\theta))=>\newline
\dfrac{\cos(5\theta)}{\cos(\theta)}=12\sin^4(\theta)-11\sin^2(\theta)+1"